OpenStudy (anonymous):

a small manufacturers weekly profit in peso is given by P=3xsquared + 70x find the number of items x that must be produced to realize a profit of P4,200 EQUATIONS : 4,200=-3xsquared -70x +3x squared -70x + 4,200 =0 .. QUESTION .. what will be the answer .. plz answer , i need it on my retrice !

OpenStudy (callisto):

Simply looking for answer is against our Code of Conduct ( http://openstudy.com/code-of-conduct). We are here to help you understand how to do it, but not help you do your homework. As, for your question, is the equation ''4,200=-3xsquared -70x'' given in the question or you did it yourself?

OpenStudy (anonymous):

nope!

OpenStudy (callisto):

That means it's not given in the question?

OpenStudy (anonymous):

was not found ?

OpenStudy (anonymous):

i dont know ?

OpenStudy (callisto):

Is the following your workings? EQUATIONS : 4,200=-3xsquared -70x +3x squared -70x + 4,200 =0

OpenStudy (callisto):

It's quite strange to me... In your question: \(P=3x^2+ 70x\) find the number of items x that must be produced to realize a profit of 4,200 Put P = 4200 into the equation. \(4200=3x^2+ 70x\) \(3x^2+ 70x-4200 = 0\) Solve it using quadratic formula \(x=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\) Identify a, b and c for your question. Can you do it?

OpenStudy (anonymous):

no .. im not good in mth!

OpenStudy (callisto):

Hmm.. the general quadratic equation is \(ax^2 + bx +c = 0\) Now you have \(3x^2 + 70x - 4200 = 0\) Can you compare the two now?

OpenStudy (anonymous):

no

OpenStudy (callisto):

What is the coefficient of \(x^2\) ?

OpenStudy (anonymous):

x ?

OpenStudy (callisto):

No. Coefficient of \(x^2\) is the number (or a variable, sometimes) in front of \(x^2\).